- 343: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Sun 15-Aug-2004 17:50:00 GMT, LMM POWER DOWN AND PROJECT TERMINATION
THE LMM SITE WAS POWERED DOWN SHORTLY AFTER THE 18Z SOUNDING ON 15 AUG 04.
ALL SUFRACE AND UPPER AIR DATA CEASE AFTER THAT TIME. THIS MARKS THE END
OF THE NAME DATA COLLECTION AT THIS SITE.
- 336: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Sat 14-Aug-2004 18:14:55 GMT, NOTES TO TEAR DOWN CREW
The notes to the tear down crew have been posted on the bulletin board
of the trailer. The security was informed that NCAR will vacate the
premises on or about 8-20.
- 329: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Fri 13-Aug-2004 20:11:39 GMT, LOSS OF MOTOSAT
The motosat went down around the 16Z sounding on 8-13. We have a roving crew
in LMM that will hopefully get us back on line by this evening. In any case,
soundings are being taken and the data are being stored. We have instructed
roving crew/night shift to send data to GTS as soon as motosat is restored.
- 290: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Mon 09-Aug-2004 00:11:26 GMT, Satellite Phone Problems
The satellite phone at LMM does not seem to be working properly. Over the past couple of days, in looking at the log, calls were limited to less than a
minute. In using it myself, I was cut-off after about 30 seconds. Whether the other calls were cut-off as well, I do not know. The only other documentation
of this possible problem was in the Morning Crew's August 8th Daily Log. As
the satellite phone is one of our primary modes of communication, especially
since cell-phone coverage can be weak at times, it is important that this issue be resolved. We will attempt to use the satellite phone again, as this issue
may just be fluke, but any insight or troubleshooting would be appreciated.
You can email or call us on our cell at 668-102-1974 with any suggestions.
- 289: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Sun 08-Aug-2004 23:18:23 GMT, SONDE TRANSPORT TO LORETO--SPECIFICS
If sonde transport to Loreto is needed, we may send them via Aeropacifico
from Los Mochis. The cost would be $12.00 MP per kilo. The flight schedule
is as follows:
Mon, Wed., Fri.
LMM Depart 8:30
LTO Arrive 14:30
The contact information for the Aeropacifico agent at the airport is (a
hardcopy of the info is in the trailer but I include it here in case it
7:00am-9am Aeropacifico Counter
2:00pm-10pm Aerocalifornia Counter
The package may be dropped off any day of the week, even if there is not a
flight that day.
Since I will be in Mazatlan Tuesday-Thursday of this week, it would be MOST
helpful if we could have a final decision by Monday afternoon (8-9) because
I am the only person in the LMM groupo that can speak Spanish and Sr. Romero
does not speak English.
- 272: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Fri 06-Aug-2004 19:59:21 GMT, SONDE TRANSPORT TO LORETO
While the roving crew was here earlier this week, Lou mentioned the possibility
that sondes may need to be transported via air cargo to Loreto. To the
roving crew or Mike, please provide the number of sondes you need to transport
and when they need to be sent by next Monday morning (8-9) if you need to do
this. See 8-6 morning shift log for more details.
- 224: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Sun 01-Aug-2004 01:15:05 GMT, REMAINING SONDES AT LMM
Mike requested that we provide the information on the remaining sondes at LMM.
Here are all our sonde ID nos. as of this evening, not including the 4Z and 8Z
launches for 8-1. Hopefully this information will be of benefit to ATD and
the roving crew in helping us solve the problem with the 408 sondes.
Sondes with 2 coefs: these don't seem to have any apparent problems. We
have been entering in coefs manually and will try the file Mike sent us that
has the coef info. Thanks, Mike for the file
Sondes with 4 coef: These have been problematic. Three good launches, two bad
ones, and a few bad sondes that weren't launched because the surface data
was grossly inaccurate
The 4 sondes that HAVE NOT had an attempted launch are:
The following 408 sondes had bad surface data and were not launched:
For the lauched "4" sonde information, please refer to the sounding logs.
These sondes remain on site and have been placed in a "problem sonde" box
so that the roving crew may inspect them.
The big question here is: how many of these unopened 408 sondes are good or
bad? As I see it, the only way to be absolutely sure of how many good sondes
we have left here would be to test the remaining unopened ones one by one.
I will leave it to ATD and the roving crew to instruct us on how to proceed.
A member of the roving crew should arrive here this week and hopefully
they will be able to investigate this problem and provide a possible solution.
We are OKAY with our sondes until then.
- 207: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Thu 29-Jul-2004 07:15:19 GMT, ALL SONDES NOW OK
The problems that occurred on the two previous shifts with sonde coeffiecients
do not seem to be an issue anymore. The sondes beginning with "408" (2 boxes)
are working without problems, and the sondes beginning with "2" work with the
manual entry of the sonde coeffiecients. Problems must have just been a streak
of bad sondes. No more sonde shortage problems are expected for the time being.
Sondes Left: 87
Coeff. 3: 4
Coeff. 4: 54
Coeff. 2: 29 (The coefficients for these still need to be entered
manually because there is not a file for them.)
- 201: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Wed 28-Jul-2004 19:01:21 GMT, Failed Attempt to manually enter coefficients for "4" soundes..
As per Steve and Roving crew request, Brad and I attempted to manually enter
coefficient valuesfor a "4" sounde in GLASS system for 20Z sounding.
To no avail, identical results were produced by GLASS where Temp, Pressure, and Moisture measurements were completely inaccurate by same factor as yesterday
(e.g. Pres 1044.7 mb).
We speculate that these coefficient values ARE correct for the "4" soundes,
either by manual entry or from floppy file upload. Given the fact that
coefficient entry always passes GLASS "synchronization" and "initialization"
phases, we believe that the problem remains exclusive to the sounde itself,
in that its just simply faulty and malfunctional.
Since we tested this method only once on our last shift launch, Brad and I will
attempt to manually enter coefficients again tomorrow on a different "4" sounde and "2" sounde.
We also ask that the same manual entry be done for the "2" soundes first
upon night shifts arrival.
IF neither of these group of soundes work...ISS 4 is up shit's creek.
- 197: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Wed 28-Jul-2004 08:42:50 GMT, Need sonde coefficient info ASAP
PLEASE READ E-MAIL SENT TO MIKE SUSEDIK AND DAILY LOG REGARDING THE SITUATION
WITH SONDE COEFFICIENTS. THIS PROBLEM MUST BE RESOLVED WITHIN THE NEXT 24h or
LMM IS OUT OF SONDES AS OF THE CURRENT IOP.
- 139: IMPORTANT IMPORTANT, Site ISS4 (Los Mochis), Tue 20-Jul-2004 10:38:16 GMT, UPS malfunction ISS 4 Los Mochis
One of the UPSs appears to be malfunctioning as a result of a power down
exercise this morning. We pulled out the power strips from this UPS and
set those in the remaining 2 functioning UPSs. This will act as a temporary
replacement until the malfunctioning UPS can be replaced by crew members,
as we are unsure of the exact power sources that were connected to this one
and the electrical load on the system.
The power down was executed because there were three missed launches for the
8Z sounding due to computer malfunctions.